OpenStudy (anonymous):

Medal Simplify. (− 5/6) + 5/36 ÷ (− 5/6)

OpenStudy (campbell_st):

well you need to use order of operations so do the division 1st... when dividing fractions flip the 2nd fraction and multiply so you have \[\frac{5}{36} \div \frac{-5}{6} = \frac{5}{36} \times \frac{-6}{5}\] when you get the answer, just add it to -5/6

OpenStudy (anonymous):

so the answer is -5/6 or -6/5

OpenStudy (campbell_st):

no...